The following information pertains to Tiller Co.:

Sales $ 710,000
Variable Costs 213,000
Fixed Costs 28,700


What is Tiller's break-even point in sales dollars? (CPA adapted) (Round intermediate calculation to 2 decimal places)

Multiple Choice

  • $241,700.

  • $41,000.

  • $213,000.

  • $28,700.
Answer #1

Contribution margin=Sales-Variable cost

=(710,000-213,000)=$497,000

Contribution margin ratio=Contribution margin/Sales

497,000/710,000=0.7

Breakeven=Fixed expenses/Contribution margin ratio

=(28700/0.7)

=$41000